# von Neumann Orbit Equivalence

### v2

@@ 3.3. Relationship to von Neumann equivalence
Suppose $(\mathcal{H},Q,\xi)$ is a triple as in condition (1) of Theorem 3.1. As in the proof of (1) implies (2) in Theorem 3.1, let $U:\mathcal{H}\to L^{2}(A\,\overline{\otimes}\,Q)$ be the unitary such that $U(x\xi)=\hat{x}$ for all $x\in A\,\overline{\otimes}\,Q$, and let $\phi:B\to A\,\overline{\otimes}\,Q$ be the *-homomorphism obtained therein. Let $\mathcal{M}=Q^{\prime}\cap\mathcal{B}(\mathcal{H})=\mathcal{B}(L^{2}(A))\,\overline{\otimes}\,Q^{\mathrm{op}}$. We will show that $\mathcal{M}$ is a von Neumann coupling between $A$ and $B$. It is clear that the inclusion $A\subset\mathcal{M}$ has a finite fundamental domain. We recall that the argument used in defining $\phi$, shows that we have an inclusion $B^{\mathrm{op}}\subset\mathcal{M}$ and moreover, since the left $A$- and right $B$-actions on $\mathcal{H}$ commute, we have that $B^{\mathrm{op}}\subset A^{\prime}\cap\mathcal{M}$. Thus, it only remains to show that the inclusion $B^{\mathrm{op}}\subset\mathcal{M}$ has a finite fundamental domain. To this end, note that, we can also view $\mathcal{H}$ as an $A-B\,\overline{\otimes}\,Q^{\mathrm{op}}$-bimoudle, and $\xi$ is tracial and cyclic for the right $B\,\overline{\otimes}\,Q^{\mathrm{op}}$-module structure. Thus, by the same construction as above, we get an inclusion $B^{\mathrm{op}}\subset{Q^{\mathrm{op}}}^{\prime}\cap\mathcal{B}(\mathcal{H})=\mathcal{B}(L^{2}(B))\,\overline{\otimes}\,Q$. Since $\mathcal{M}=Q\cap\mathcal{B}(\mathcal{H})={Q^{\mathrm{op}}}^{\prime}\cap\mathcal{B}(H)$, we get that the inclusion $B^{\mathrm{op}}\subset\mathcal{M}$ admits a finite fundamental domain and hence $A\sim_{\mathrm{vNE}}B$. ∎
We conclude this section by proving Theorem 1.8, which we recall below.
@@
Proof.
Let $(Q_{v},\phi_{v})$ be a vNOE-coupling between $(A,\tau_{A_{v}})$ and $(B_{v},\tau_{B_{v}})$, where $\phi_{v}:A_{v}\to B_{v}\,\overline{\otimes}\,Q_{v}$ is a normal unital *-homomorphism satisfying
here, ${\mathbb{E}}_{v}:B_{v}\,\overline{\otimes}\,Q_{v}\to Q_{v}$ is the normal conditional expectation. Define $(Q,\tau_{Q})$ to be the tensor product ¯$\,\overline{\otimes}\,_{v\in V}(Q_{v},\tau_{Q_{v}})$. We now define a *-homomorphism $\phi:A\to B\,\overline{\otimes}\,Q$ and show that it has the desired properties. To this end, first note that if $v$ and $w$ are two distinct vertices connected by an edge, then $B_{v}$ commutes with $B_{w}$ and $Q_{v}$ commutes with $Q_{w}$ as $v\neq w$. Thus, $\phi_{v}(A_{v})$ and $\phi_{w}(A_{w})$ commute inside $B\,\overline{\otimes}\,Q$, and therefore, we can define $\phi$ on the algebraic graph product (i.e., the universal unital *-algebra generated by the $\{A_{v}\}_{v\in V}$ subject to the relation that $A_{v}$ commutes with $A_{w}$ whenever $v$ and $w$ are connected by an edge). In particular, note that $\phi(a)=\phi_{v}(a)$, whenever $a\in A_{v}$. Next, we show that $\phi$ extends to a normal unital *-homomorphism on $A$ and satisfies ${\mathbb{E}}_{Q}(\phi(a))=\tau_{A}(a)$ for all $a\in A$, where ${\mathbb{E}}_{Q}:B\,\overline{\otimes}\,Q\to Q$ is the normal conditional expectation. To see this, we note that for any reduced word $v=v_{1}\cdots v_{n}$ of vertices, and elements $a_{i}\in A_{v_{i}}$ of trace zero, $1\leq i\leq n$, we have that
${\mathbb{E}}_{Q}(\phi_{1}(a_{1})\cdots\phi_{n}(a_{n}))=0.$
Indeed, since ${\mathbb{E}}_{v_{i}}(\phi_{v_{i}}(a_{v_{i}}))=\tau_{A_{v_{i}}}(a_{v_{i}})=0,~{}1\leq i\leq n$, so, $\phi_{v_{i}}(a_{v_{i}})$ can be approximated by linear combinations of elements of the form $b_{v_{i}}\otimes q_{v_{i}}$, where $b_{v_{i}}\in B_{v_{i}},q_{v_{i}}\in Q_{v_{i}}$, and $\tau_{B_{v_{i}}}(b_{v_{i}})=0,~{}1\leq i\leq n$. Thus, ${\mathbb{E}}_{Q}(\phi_{1}(a_{1})\cdots\phi_{n}(a_{n}))=0$ and hence it follows that the map $\phi$, defined on the algebraic graph product, is trace-preserving. Therefore, by [CF17, Proposition 3.22], $\phi$ extends to a well-defined normal unital *-homomorphism on $A$ and satisfies ${\mathbb{E}}_{Q}(\phi(a))=\tau_{A}(a)$ for all $a\in A$. Finally, to show that $\overline{\operatorname{Span}\{\phi(a)x:a\in A,x\in Q\}}^{\|\cdot\|_{2}}=L^{2}(B\,\overline{\otimes}\,Q)$, it suffices to show that
for any reduced word $v=v_{1}\cdots v_{n}$ of vertices and for all $n$. We proceed by induction. It is straightforward to see that for any vertex $v$, $\overline{\operatorname{Span}(\phi_{v}(A_{v})(1\otimes Q))}^{\|\cdot\|_{2}}=L^{2}(B_{v}\,\overline{\otimes}\,Q)=\overline{\operatorname{Span}(B_{v}(1\otimes Q))}^{\|\cdot\|_{2}}$. Suppose that (*3.3) holds for any reduced word $v_{1}\cdots v_{n}$ of length $n$ and let $v_{1}\cdots v_{n}v_{n+1}$ be any reduced word of length $n+1$. Then,
$\overline{\operatorname{Span}(\phi_{v_{1}}(A_{v_{1}})\cdots\phi_{n+1}(A_{n+1})(1\otimes Q))}^{\|\cdot\|_{2}}$ $=\overline{\operatorname{Span}(\phi_{v_{1}}(A_{v_{1}})B_{v_{2}}\cdots B_{v_{n+1}}(1\otimes Q))}^{\|\cdot\|_{2}}$
Thus, $(A,\tau_{A})\sim_{\mathrm{vNOE}}(B,\tau_{B})$. ∎
4. Towards an analogue of Singer’s Theorem
Let $\Gamma$ be a countable discrete group and $(M,\tau)$ be a finite von Neumann algebra. A 1-cocycle for a trace-preserving action $\Gamma{\,\curvearrowright\,}^{\alpha}(M,\tau)$ is a map $w:\Gamma\to\mathcal{U}(M)$ that satisfies the following cocycle identity:
@@ 4. Towards an analogue of Singer’s Theorem
Let $\Gamma{\,\curvearrowright\,}^{\sigma}\mathcal{M}$ and $\Lambda{\,\curvearrowright\,}^{\alpha}\mathcal{M}$ be commuting, trace-preserving actions of countable discrete groups $\Gamma$ and $\Lambda$ on a semi-finite von Neumann algebra $\mathcal{M}$ with a faithful normal semi-finite trace $\mathrm{Tr}$. Let $p\in\mathcal{M}$ be a finite-trace projection which is a common fundamental domain for both $\Gamma$- and $\Lambda$-actions, that is, $\{\sigma_{\gamma}(p)\}_{\gamma\in\Gamma}$ are mutually orthogonal and $\sum_{\gamma\in\Gamma}\sigma_{\gamma}(p)=1$ (SOT), and similarly, $\{\alpha_{\lambda}(p)\}_{\lambda\in\Lambda}$ are mutually orthogonal and $\sum_{\lambda\in\Lambda}\alpha_{\lambda}(p)=1$ (SOT). From [IPR24, Propostion 4.2], there exists a unitary $\mathcal{F}_{p}:\ell^{2}\Gamma\,\overline{\otimes}\,L^{2}(\mathcal{M}^{\Gamma},\mathrm{Tr})\to L^{2}(\mathcal{M},\mathrm{Tr})$ such that $\mathcal{F}_{p}(\delta_{\gamma}\otimes x)=\sigma_{\gamma^{-1}}(p)x$ for all $\gamma\in\Gamma,x\in\mathcal{M}^{\Gamma}$. Furthermore, from [IPR24, Proposition 4.3], there is a trace-preserving isomorphism $\Delta_{p}^{\Gamma}:\mathcal{M}\rtimes\Gamma\to\mathcal{B}(\ell^{2}\Gamma)\,\overline{\otimes}\,\mathcal{M}^{\Gamma}$ such that for $\gamma\in\Gamma$ and $x\in\mathcal{M}$,
@@
$(\mathcal{M}\rtimes\Gamma)\rtimes_{\alpha\rtimes\operatorname{id}_{\Gamma}}\Lambda\cong(\mathcal{B}(\ell^{2}\Gamma)\,\overline{\otimes}\,\mathcal{M}^{\Gamma})\rtimes_{\tilde{\alpha}}\Lambda.$
Now, for any $x\in\mathcal{M},\gamma\in\Gamma$, and $\lambda\in\Lambda$, we have
$\Delta_{p}^{\Gamma}\circ(\alpha_{\lambda}\rtimes\operatorname{id}_{\Gamma})(xu_{\gamma})$ $=\Delta_{p}^{\Gamma}(\alpha_{\lambda}(x))(\rho_{\gamma}\otimes 1)$
@@
$=(\operatorname{id}\otimes\alpha_{\lambda})(v_{\lambda}\Delta_{p}^{\Gamma}(xu_{g})v_{\lambda}^{*}),$
where $v_{\lambda}=\mathcal{F}^{*}_{\alpha_{\lambda^{-1}}(p)}\mathcal{F}_{p}$. The last equality follows from the fact that $v_{\lambda}\in\mathcal{U}(L\Gamma\,\overline{\otimes}\,\mathcal{M}^{\Gamma})$ (see [IPR24, Proposition, 4.4]), and hence commutes with $\rho_{\gamma}\otimes 1$. Therefore, if we let $w_{\lambda}=(\operatorname{id}\otimes\alpha_{\lambda})(v_{\lambda})\in\mathcal{U}(L\Gamma\,\overline{\otimes}\,\mathcal{M}^{\Gamma})$, then we have that
@@
Proof.
First note that, for any $x\in\mathfrak{n}_{\mathrm{Tr}}$, it is straightforward to verify that
$\mathcal{F}_{p}^{*}(x)=\sum_{\gamma\in\gamma}\delta_{\gamma}\otimes x_{\gamma},$
@@
$x_{\gamma}=\sum_{b\in\Gamma}\sigma_{b\gamma^{-1}}(p)\sigma_{b}(x).$
Therefore, for any $a\in\Gamma$ and $x\in\mathcal{M}^{\Gamma}$, we have
@@
$=\sum_{\gamma\in\Gamma}\delta_{\gamma}\otimes\left(\sum_{b\in\Gamma}\sigma_{b\gamma^{-1}}(\alpha_{\lambda^{-1}}(p))\sigma_{b}(\sigma_{a^{-1}}(p)x\right)$
@@
It now follows from the above claim that the actions $\tilde{\alpha}$ and $\operatorname{id}\otimes\alpha$ of $\Lambda$ on $\mathcal{B}(\ell^{2}\Gamma)\,\overline{\otimes}\,\mathcal{M}^{\Gamma}$ are cocycle conjugate, and hence we get an isomorphism of the crossed products
@@
$(\mathcal{B}(\ell^{2}\Gamma)\,\overline{\otimes}\,\mathcal{M}^{\Gamma})\rtimes_{\tilde{\alpha}}\Lambda\underset{\cong}{\overset{\Psi^{\Gamma}}{\longrightarrow}}(\mathcal{B}(\ell^{2}\Gamma)\,\overline{\otimes}\,\mathcal{M}^{\Gamma})\rtimes_{\operatorname{id}\otimes\alpha}\Lambda\cong\mathcal{B}(\ell^{2}\Gamma)\,\overline{\otimes}\,(\mathcal{M}^{\Gamma}\rtimes_{\alpha}\Lambda).$
@@
Similarly, starting with the isomorphism $\Delta_{p}^{\Lambda}:\mathcal{M}\rtimes\Lambda\to\mathcal{B}(\ell^{2}\Lambda)\,\overline{\otimes}\,\mathcal{M}^{\Lambda}$, and performing the above analysis yields the following isomorphism of the crossed products
@@
$(\mathcal{B}(\ell^{2}\Lambda)\,\overline{\otimes}\,\mathcal{M}^{\Lambda})\rtimes_{\tilde{\sigma}}\Gamma\underset{\cong}{\overset{\Psi^{\Lambda}}{\longrightarrow}}(\mathcal{B}(\ell^{2}\Lambda)\,\overline{\otimes}\,\mathcal{M}^{\Lambda})\rtimes_{\operatorname{id}\otimes\sigma}\Gamma\cong\mathcal{B}(\ell^{2}\Lambda)\,\overline{\otimes}\,(\mathcal{M}^{\Lambda}\rtimes_{\sigma}\Gamma).$
Thus, there exists an isomorphism $\Phi:\mathcal{B}(\ell^{2}\Gamma)\,\overline{\otimes}\,(\mathcal{M}^{\Gamma}\rtimes_{\alpha}\Lambda)\to\mathcal{B}(\ell^{2}\Lambda)\,\overline{\otimes}\,(\mathcal{M}^{\Lambda}\rtimes_{\sigma}\Gamma)$ making the following diagram commutative.
[…]
If we let $\omega_{e_{\Gamma},e_{\Gamma}}\in\mathcal{B}(\ell^{2}\Gamma)$ (resp. $\omega_{e_{\Lambda},e_{\Lambda}}\in\mathcal{B}(\ell^{2}\Lambda)$) denote the orthogonal projection onto $\mathbb{C}\delta_{e_{\Gamma}}$ (resp. $\mathbb{C}\delta_{e_{\Lambda}}$), then we note that
$\Phi(\omega_{e_{\Gamma},e_{\Gamma}}\otimes 1)=\Psi^{\Lambda}(\theta(\triangle_{p}^{\Gamma}(p)))=\omega_{e_{\Lambda},e_{\Lambda}}\otimes 1,$
and therefore, we have

### v3

@@ 3.3. Relationship to von Neumann equivalence
Suppose $(\mathcal{H},Q,\xi)$ is a triple as in condition (1) of Theorem 3.1. As in the proof of (1) implies (2) in Theorem 3.1, let $U:\mathcal{H}\to L^{2}(A\,\overline{\otimes}\,Q)$ be the unitary such that $U(x\xi)=\hat{x}$ for all $x\in A\,\overline{\otimes}\,Q$, and let $\phi:B\to A\,\overline{\otimes}\,Q$ be the *-homomorphism obtained therein. Let $\mathcal{M}=Q^{\prime}\cap\mathcal{B}(\mathcal{H})=\mathcal{B}(L^{2}(A))\,\overline{\otimes}\,Q^{\mathrm{op}}$. We will show that $\mathcal{M}$ is a von Neumann coupling between $A$ and $B$. It is clear that the inclusion $A\subset\mathcal{M}$ has a finite fundamental domain. We recall that the argument used in defining $\phi$, shows that we have an inclusion $B^{\mathrm{op}}\subset\mathcal{M}$ and moreover, since the left $A$- and right $B$-actions on $\mathcal{H}$ commute, we have that $B^{\mathrm{op}}\subset A^{\prime}\cap\mathcal{M}$. Thus, it only remains to show that the inclusion $B^{\mathrm{op}}\subset\mathcal{M}$ has a finite fundamental domain. To this end, note that, we can also view $\mathcal{H}$ as an $A-B\,\overline{\otimes}\,Q^{\mathrm{op}}$-bimodule, and $\xi$ is tracial and cyclic for the right $B\,\overline{\otimes}\,Q^{\mathrm{op}}$-module structure. Thus, by the same construction as above, we get an inclusion $B^{\mathrm{op}}\subset{Q^{\mathrm{op}}}^{\prime}\cap\mathcal{B}(\mathcal{H})=\mathcal{B}(L^{2}(B))\,\overline{\otimes}\,Q$. Since $\mathcal{M}=Q^{\prime}\cap\mathcal{B}(\mathcal{H})={Q^{\mathrm{op}}}^{\prime}\cap\mathcal{B}(\mathcal{H})$, we get that the inclusion $B^{\mathrm{op}}\subset\mathcal{M}$ admits a finite fundamental domain and hence $A\sim_{\mathrm{vNE}}B$. ∎
4. Towards an analogue of Singer’s Theorem
The main goal of this section is to prove the following theorem.
@@ 4. Towards an analogue of Singer’s Theorem
Before proving the theorem, we first set up some notations and recall a few facts. Let $\Gamma$ be a countable discrete group and $(M,\tau)$ be a tracial von Neumann algebra. A 1-cocycle for a trace-preserving action $\Gamma{\,\curvearrowright\,}^{\alpha}(M,\tau)$ is a map $w:\Gamma\to\mathcal{U}(M)$ that satisfies the following cocycle identity:
@@
Proof of Theorem 4.1.
Let $\Gamma{\,\curvearrowright\,}^{\sigma}\mathcal{M}$ and $\Lambda{\,\curvearrowright\,}^{\alpha}\mathcal{M}$ be commuting, trace-preserving actions of countable discrete groups $\Gamma$ and $\Lambda$ on a semi-finite von Neumann algebra $\mathcal{M}$ with a faithful normal semi-finite trace $\mathrm{Tr}$. Let $p\in\mathcal{M}$ be a finite-trace projection which is a common fundamental domain for both $\Gamma$- and $\Lambda$-actions, that is, $\{\sigma_{\gamma}(p)\}_{\gamma\in\Gamma}$ are mutually orthogonal and $\sum_{\gamma\in\Gamma}\sigma_{\gamma}(p)=1$ (SOT), and similarly, $\{\alpha_{\lambda}(p)\}_{\lambda\in\Lambda}$ are mutually orthogonal and $\sum_{\lambda\in\Lambda}\alpha_{\lambda}(p)=1$ (SOT). From [IPR24, Propostion 4.2], there exists a unitary $\mathcal{F}_{p}:\ell^{2}\Gamma\,\overline{\otimes}\,L^{2}(\mathcal{M}^{\Gamma},\tau)\to L^{2}(\mathcal{M},\mathrm{Tr})$ such that
$\mathcal{F}_{p}(\delta_{\gamma}\otimes x)=\sigma_{\gamma^{-1}}(p)x\qquad\text{for all }\gamma\in\Gamma,x\in\mathcal{M}^{\Gamma},$
where the trace $\tau$ on $\mathcal{M}^{\Gamma}$ is given by $\tau(x)=\mathrm{Tr}(pxp),~x\in\mathcal{M}^{\Gamma}$ [^2] Note that the formula for $\mathcal{F}_{p}$ in [IPR24, Proposition 4.2] requires $x\in\mathfrak{n}_{\tau}\subset\mathcal{M}^{\Gamma}$, where $\tau$ is the semi-finite trace on $\mathcal{M}^{\Gamma}$ given by $\tau(x)=\mathrm{Tr}(pxp)$. However, when $\mathrm{Tr}(p)<\infty$, $\tau$ is a finite trace and thus we can have $x\in\mathcal{M}^{\Gamma}$ in the formula for $\mathcal{F}_{p}$.. Furthermore, from [IPR24, Proposition 4.3], there is a trace-preserving isomorphism $\Delta_{p}^{\Gamma}:\mathcal{M}\rtimes\Gamma\to\mathcal{B}(\ell^{2}\Gamma)\,\overline{\otimes}\,\mathcal{M}^{\Gamma}$ such that for $\gamma\in\Gamma$ and $x\in\mathcal{M}$,
@@
$\mathcal{M}\rtimes(\Gamma\times\Lambda)=(\mathcal{M}\rtimes\Gamma)\rtimes_{\alpha\rtimes\operatorname{id}_{\Gamma}}\Lambda\underset{\cong}{\overset{\Delta_{p}^{\Gamma}\rtimes\operatorname{id}_{\Gamma}}{\longrightarrow}}(\mathcal{B}(\ell^{2}\Gamma)\,\overline{\otimes}\,\mathcal{M}^{\Gamma})\rtimes_{\tilde{\alpha}}\Lambda.$
Now we note that for any $\lambda\in\Lambda$, $\alpha_{\lambda^{-1}}(p)$ is again a $\Gamma$-fundamental domain and hence we can consider the map $\Delta_{\alpha_{\lambda^{-1}}(p)}^{\Gamma}:\mathcal{M}\rtimes\Gamma\to\mathcal{B}(\ell^{2})\otimes\mathcal{M}^{\Gamma}$ defined analogously to $\Delta_{p}^{\Gamma}$. For any $x\in\mathcal{M}$ and $s,t\in\Gamma$, by a slight abuse of notation, let $\Delta_{p}^{\Gamma}(x)]_{s,t}$ be the $(s,t)$-entry in the matrix representation of $\Delta_{p}^{\Gamma}(x)$. Similarly, consider $\Delta_{\alpha_{\lambda^{-1}}(p)}^{\Gamma}(x)]_{s,t}$. We observe that
$\alpha_{\lambda}([\Delta_{\alpha_{\lambda^{-1}}(p)}^{\Gamma}(x)]_{s,t})$ $=\alpha_{\lambda}\left(\sum_{\gamma\in\Gamma}\sigma_{\gamma}(\sigma_{t^{-1}}(\alpha_{\lambda^{-1}}(p))x\sigma_{s^{-1}}(\alpha_{\lambda^{-1}}(p)))\right)$
$=\sum_{\gamma\in\Gamma}\sigma_{\gamma}(\sigma_{t^{-1}}(p)\alpha_{\lambda}(x)\sigma_{s^{-1}}(p))$
$=[\Delta_{p}^{\Gamma}(\alpha_{\lambda}(x))]_{s,t},$
whence it follows that $\Delta_{p}^{\Gamma}(\alpha_{\lambda}(x))=(\operatorname{id}\otimes\alpha_{\lambda})(\Delta_{\alpha_{\lambda^{-1}}(p)}^{\Gamma}(x))$ for all $x\in\mathcal{M}$ and $\lambda\in\Lambda$. Therefore, for any $x\in\mathcal{M},\gamma\in\Gamma$, and $\lambda\in\Lambda$, we have
$(\Delta_{p}^{\Gamma}\circ(\alpha_{\lambda}\rtimes\operatorname{id}_{\Gamma}))(xu_{\gamma})$ $=\Delta_{p}^{\Gamma}(\alpha_{\lambda}(x)u_{\gamma})$
$=\Delta_{p}^{\Gamma}(\alpha_{\lambda}(x))\Delta_{p}^{\Gamma}(u_{\gamma})$
@@
$=(\operatorname{id}\otimes\alpha_{\lambda})(v_{\lambda}\Delta_{p}^{\Gamma}(xu_{\gamma})v_{\lambda}^{*}),$
where $v_{\lambda}=\mathcal{F}^{*}_{\alpha_{\lambda^{-1}}(p)}\mathcal{F}_{p}$. The last equality follows from the fact that $v_{\lambda}\in\mathcal{U}(L\Gamma\,\overline{\otimes}\,\mathcal{M}^{\Gamma})$ (see [IPR24, Proposition 4.4]), and hence commutes with $\rho_{\gamma}\otimes 1$. Therefore, if we let $w_{\lambda}=(\operatorname{id}\otimes\alpha_{\lambda})(v_{\lambda})\in\mathcal{U}(L\Gamma\,\overline{\otimes}\,\mathcal{M}^{\Gamma})$, then we have that
@@
Proof of Claim.
First note that, for any $y\in\mathfrak{n}_{\mathrm{Tr}}\subset\mathcal{M}$, it is straightforward to verify that
$\mathcal{F}_{p}^{*}(y)=\sum_{\gamma\in\Gamma}\delta_{\gamma}\otimes y_{\gamma},$
@@
$y_{\gamma}=\sum_{b\in\Gamma}\sigma_{b\gamma^{-1}}(p)\sigma_{b}(y).$
Since $\mathrm{Tr}(p)<\infty$, so, for any $a\in\Gamma$ and $x\in\mathcal{M}^{\Gamma}$, we have that $\mathrm{Tr}(\sigma_{a^{-1}}(p)x)<\infty$ and hence $\sigma_{a^{-1}}(p)x\in\mathfrak{n}_{\mathrm{Tr}}$. Therefore, for any $a\in\Gamma$ and $x\in\mathcal{M}^{\Gamma}$, we have
@@
$=\sum_{\gamma\in\Gamma}\delta_{\gamma}\otimes\left(\sum_{b\in\Gamma}\sigma_{b\gamma^{-1}}(\alpha_{\lambda^{-1}}(p))\sigma_{b}(\sigma_{a^{-1}}(p)x)\right)$
@@
It now follows from the above claim that the actions $\tilde{\alpha}$ and $\operatorname{id}\otimes\alpha$ of $\Lambda$ on $\mathcal{B}(\ell^{2}\Gamma)\,\overline{\otimes}\,\mathcal{M}^{\Gamma}$ are cocycle conjugate, and hence we get the following isomorphisms of the crossed products
@@
$(\mathcal{B}(\ell^{2}\Gamma)\,\overline{\otimes}\,\mathcal{M}^{\Gamma})\rtimes_{\tilde{\alpha}}\Lambda\underset{\cong}{\overset{\Psi^{\Gamma}}{\longrightarrow}}(\mathcal{B}(\ell^{2}\Gamma)\,\overline{\otimes}\,\mathcal{M}^{\Gamma})\rtimes_{\operatorname{id}\otimes\alpha}\Lambda\underset{\cong}{\overset{\Psi^{\alpha}}{\longrightarrow}}\mathcal{B}(\ell^{2}\Gamma)\,\overline{\otimes}\,(\mathcal{M}^{\Gamma}\rtimes_{\alpha}\Lambda),$
where the isomorphism $\Psi^{\alpha}$ is the canonical isomorphism and satisfies
$\Psi^{\alpha}((T\otimes x)u_{\lambda})=T\otimes xu_{\lambda},\qquad T\in\mathcal{B}(\ell^{2}\Gamma),x\in\mathcal{M}^{\Gamma},\lambda\in\Lambda.$
@@
Similarly, starting with the isomorphism $\Delta_{p}^{\Lambda}:\mathcal{M}\rtimes\Lambda\to\mathcal{B}(\ell^{2}\Lambda)\,\overline{\otimes}\,\mathcal{M}^{\Lambda}$, and performing the above analysis, yields the following isomorphisms of the crossed products
@@
$\mathcal{M}\rtimes(\Gamma\times\Lambda)=(\mathcal{M}\rtimes\Lambda)\rtimes_{\sigma\rtimes\operatorname{id}_{\Lambda}}\Gamma\underset{\cong}{\overset{\Delta_{p}^{\Lambda}\rtimes\operatorname{id}_{\Lambda}}{\longrightarrow}}(\mathcal{B}(\ell^{2}\Lambda)\,\overline{\otimes}\,\mathcal{M}^{\Lambda})\rtimes_{\tilde{\sigma}}\Gamma,$